User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Note to observers: Aircon in Auscope control room appears to be out of water or something. If you turn it on, it smells bad and doesn't really heat up the room.

Hobart 26m

Disk VSN: HOB+0015/6000

Start Pos: 4105 GB

  • Exp started OK.
  • He supply press ~275 psi. Informing Brett.
  • 5:15UTC - He was supplied, bringing pressure from 265 to 300 psi (Liza).
  • Signs of RFI in channels S5, S6 (overflow) pretty much all the time (Liza)
  • Racbus lost connection with newsmerd 2 times, the pcal control had to be returned to fs manually (calu…) Caused overflow of all VCs (Liza)
  • 11:45 He supply pressure ~278 psi. Dropping at this rate it should be below 275 psi in about an hour or two. Liza will contact Brett and he will make decision (Anita)
  • A little after 14:00 UT Brett re-filled the He-supply, now is about 298 psi. (Ellen)


Disk VSN: MED-0020/2000

Start Pos: 0 GB

  • Exp started OK
  • High Tsys in channels S1-S4 pretty much all the time (120-170K) (Liza)


Disk VSN: USN-0201/2000

Start Pos: 0 GB

  • Exp started OK.
  • Monica restarted twice due to ifpic problem (Liza)
  • Monica also restarted twice due to ifpic problem - does not seem to be fixing it (Anita)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1699.txt · Last modified: 2015/08/05 14:27 by David Horsley